OpenStudy (anonymous):

When finding the product (x – 3)(x + 4), what would you add to xsquared-3x to get the final answer? A. –4x + 12 B. –3x – 12 C. 3x + 12 D. 4x – 12

OpenStudy (lncognlto):

Well, what is (x – 3)(x + 4) expanded?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

what u mean expand

OpenStudy (lncognlto):

:), Ok, use the FOIL method. Multiply the two First ones, together, then the Outside ones, then the Inside ones, and then the Last ones and then add them all together. What do you get?

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

i think D but i may have done it wrong

OpenStudy (phi):

See http://www.khanacademy.org/math/algebra/polynomials/multiplying_polynomials/v/multiplying-binomials if you want to learn how to do this

OpenStudy (anonymous):

thanks

OpenStudy (lncognlto):

Well lets see. (x – 3)(x + 4) = x^2 + 4x - 3x -12 =x^2 + x -12 ...(1) So to get from x^2 - 3x to (1), you need to add 4x - 12. So D is indeed correct!

OpenStudy (anonymous):

thanks you

OpenStudy (lncognlto):

You're welcome :)